OpenStudy (anonymous):

What is the slope of a line parallel to 4x + 2y = 5?

OpenStudy (saifoo.khan):

Slope = -2

OpenStudy (amistre64):

one of the simplest ways to find slope is to have x=0; then x=1 .... and measure the change in y values. 4(0) + 2y = 5 ; y = 5/2 5/2 + m = 1/2; m=-4/2, -2 4(1) +2y = 5 ; y = 1/2

OpenStudy (amistre64):